Find dy/dx and simply the result, if possible.(with solution)

A.     y=√x -(1/√x)

B.     y=x^2+π^2+x^π

C.     y=(sin x-1)/(cos x)

D.     y=x^2 sec x

E.      y=(1)/(e^× +2)


1
Expert's answer
2021-05-07T11:31:02-0400

A)

"Y=\sqrt{x}+\dfrac{1}{\sqrt{x}}"

"\dfrac{dy}{dx}=\dfrac{1}{2\sqrt x}+\dfrac{1}{2\sqrt[3]{x}}"


B)

"Y=\ x^{2}+\pi^{2}+x^{\pi}"

"\dfrac{dy}{dx}=2x+\pi{x}^{\pi-1}"


C)

"Y=\dfrac{sinx}{cosx-1}"

"\dfrac{dy}{dx}=\dfrac{cosx\times cosx-(sinx-1)(-sinx)}{(cosx-1)^2}"

"\dfrac{dy}{dx}=\dfrac{cos^{2}x+sin^{2}x-sinx}{{{(cosx-1}})^{2}}"


"\dfrac{dy}{dx}=\dfrac{1-sinx}{(cosx-1)^{2}}"


D)

"Y=x^{2}secx"

"\dfrac{dy}{dx}=2\times x\times secx+secx\times tanx \times x^{2}"


=xsecx(xtanx+2)

E)

"Y=\dfrac{1}{e^{x}+2}"

"\dfrac{dy}{dx}=\dfrac{e^{x}+2\times0+1\times e^{x}}{(e^x+2)^2}=\dfrac{e^{x}}{(e^x+2)^2}"